# Omni-R1: Do You Really Need Audio to Fine-Tune Your Audio LLM?

### v2

@@ Abstract
Index Terms:
Audio Large Language Models (LLMs)
@@ I Introduction
Reinforcement Learning (RL) has recently been shown to improve the reasoning capabilities of Large Language Models (LLMs) [1]. We are motivated by these advancements to improve the capabilities of Audio LLMs - models which take in audio input and text and can perform tasks such as question answering. Building on Qwen2.5-Omni [2], a State-of-the-Art (SOTA) multi-modal LLM, we introduce Omni-R1 using a fine-tuning pipeline based on the RL method Group Relative Policy Optimization (GRPO) [3]. Using a simple prompt, our model forgoes complex chain-of-thought or structured reasoning outputs and instead directly outputs answer choices. We use the recent MMAU [4] and MMAR [5] benchmarks to test our models. Fine-tuning on the audio and human-annotated questions from the AVQA [6] dataset boosts Qwen2.5-Omni’s average accuracy on MMAU Test-mini from 65.9% to 68.6%, and on the Test-full split from 68.4% to 70.8%.
We also propose to automatically generate audio question answering datasets. We prompt ChatGPT with several audio captions from an audio LLM to generate questions and answer choices for the 40k audios in AVQA and 182k audios in VGGSound. Scaling up the training data with our automatically generated questions results in further gains, achieving new SOTA of 71.3% on Test-mini and 71.2% on Test full. We obtained similar gains on the MMAR [5] benchmark leading to SOTA results.
@@
We plan to release our code, models, and datasets publicly.
@@ II Related Work
Several recent large-scale audio question answering datasets were proposed to test audio LLMs, including MMAU [4] and MMAR [5]. The MMAU benchmark [4] is a fixed answer choice dataset providing one correct answer and three incorrect answers for each question. It contains questions about sounds, speech, and music of varying difficulty. Some of the questions require external world-based knowledge that is not provided in the questions themselves. MMAR [5] is a more recent benchmark designed to elicit deep reasoning from audio LLMs. It includes questions about sounds, speech, and music, and also mixtures of them (for example, audio natively containing both sounds and speech). The question and answer format is the same as MMAU.
GRPO is a Reinforcement Learning (RL) method proposed for instilling better reasoning capabilities in LLMs [3, 1]. Towards better Audio LLMs, R1-AQA [7] proposes to use GRPO by fine-tuning Qwen2-Audio [8] on the AVQA [6] dataset of audio question and answers. This method achieved the previous SOTA on MMAU, and our approach is inspired by them. Using GRPO, we fine-tune Qwen2.5-Omni-7B [2], a recent multi-modal LLM proposed to handle both audio and video inputs, on AVQA. Doing so, we were able to achieve the new SOTA on MMAU and also demonstrate strong performance on the MMAR benchmark, achieving SOTA results among open-source models and competitive performance to significantly bigger, close-ended models. Further, we propose to automatically generate audio question and answer training data, which leads to even better performance.
SARI [9] is a concurrent method which fine-tunes Qwen2.5-Omni using RL. However, their setup is more complex than ours: they fine-tune the model with a schedule of Supervised Fine-Tuning (SFT) and RL, and use both structured reasoning and unstructured reasoning. In contrast, we have a much simpler pipeline: we only fine-tune the model with RL, and we don’t use any explicit reasoning.
Finally, we investigated how GRPO improves the Audio LLMs’ performance by testing them with text-only inputs. While recent work already tested audio LLMs with text-only inputs[4, 10], we go a step further and fine-tune the models without audio (just text). We made a surprising discovery that fine-tuning the models with text-only inputs could work almost as well as fine-tuning with audio. It explains that much of the performance improvement from GRPO is due to improving the text-only reasoning capabilities.
@@ III-A Audio LLM and Prompt
We propose Omni-R1 which fine-tunes Qwen2.5-Omni-7B [2] using GRPO. During fine-tuning and inference, we use the prompt: <question> Please choose the answer from the following options: <choices>. We chose this prompt since it is simple and the model learns to directly output one of the answer choices. This is especially useful since we have limited GPU memory. Thanks to this simple prompt, we were able to perform full-finetuning on GPUs with only 48GB GPU memory, in contrast to recent works which perform fine-tuning on GPUs with 80GB GPU memory [7]. Although recent works [11, 7, 9] propose more complex prompts with reasoning, our simple prompt was sufficient for our models to achieve the new SOTA on MMAU and MMAR.
@@ III-B Reinforcement Learning
Our training uses Group Relative Policy Optimization (GRPO) [7, 3], an adaptation of Proximal Policy Optimization (PPO) [12]. GRPO is designed to mitigate the substantial memory overhead and the challenges of training an accurate per-token value function, common in standard PPO implementations for LLMs. GRPO achieves this by eliminating the explicit value function. Instead, it derives advantage estimates by comparing outputs within a group generated for the same input prompt, effectively using the average reward of sampled responses as a baseline.
@@
—– Human [4] 86.3 78.2 82.2 82.2 - - - -
Audio Flamingo 2 Direct Inference [13] 61.6 74.0 30.9 55.5 65.1 72.9 40.3 59.4
Phi-4 Multi-Modal Direct Inference [14] 61.0 52.9 52.8 55.6 - - - -
GPT-4o Voice Model Direct Inference [15] 63.4 60.8 53.2 59.1 - - - -
Qwen2-Audio-7B-Instruct Direct Inference [8] 55.0 51.0 42.0 49.2 45.9 53.3 45.9 52.5
Qwen2-Audio-7B-Instruct Reproduced [8] 61.6 54.5 42.0 52.7 - - - -
Qwen2-Audio-7B-Instruct Zero-Shot-CoT [16] 61.9 56.3 55.3 57.8 - - - -
Qwen2-Audio-7B-Instruct CoTA [11] 60.1 64.3 60.7 61.7 - - - -
Qwen2-Audio-7B-Instruct R1-AQA [7] 68.8 64.3 63.7 65.6 69.8 61.4 62.7 64.4
Kimi-Audio Direct Inference [17] 73.3 61.7 60.7 65.2 - - - -
Qwen2.5-Omni-7B Direct Inference [2] 67.9 69.2 59.8 65.6 - - - -
@@
Qwen2.5-Omni-7B SARI [9] 72.8 67.2 61.3 67.1 - - - -
@@
$\footnotesize\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{\mathcal{D}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{\{}\min\Bigg{[}\rho_{i,t}\hat{A}_{i,t},\\ &\quad\left.\operatorname{clip}\left(\rho_{i,t},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta{\mathbb{D}_{KL}}_{i,t}\left[\pi_{\theta}||\pi_{ref}\right]\Bigg{\}}\Bigg{]},\end{split}$ (3)
where $\rho_{i,t}=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}$ is the probability ratio between the current policy $\pi_{\theta}$ and the policy $\pi_{\theta_{old}}$ used for sampling and $\mathcal{D}=[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)]$ denote the sampling distribution over questions and their corresponding outputs. The $\operatorname{clip}(x,min\_val,max\_val)$ function limits the value of $x$ to be within the range $[min\_val,max\_val]$. The hyperparameter $\epsilon$ defines the PPO clipping range. The per-token KL divergence ${\mathbb{D}_{KL}}_{i,t}\left[\pi_{\theta}||\pi_{ref}\right]$ between the current policy $\pi_{\theta}$ and a reference policy $\pi_{ref}$ is weighted by the hyperparameter $\beta$:
@@ III-C Automatic Q/A Dataset Generation with ChatGPT
We propose to generate audio question answer datasets automatically. Given an audio dataset, we first acquire audio-based text captions from an audio LLM. Specifically, we use the captions generated by Qwen-2 Audio [8] in the AudioSetCaps dataset [18], which is a compilation of captions generated for many open-source audio datasets. We use multiple captions which describe the overall acoustic scene, the speech present in the audio, and the music present in the audio. Next, we use the captions and prompt ChatGPT to generate questions and answers using the provided captions. We instruct the model to generate a correct answer and three other answers which are similar but clearly incorrect (prompt shown in Figure 1). We reviewed the generated questions and answers and found the quality to be reasonable, however, we noticed that the original captions feature hallucinations especially about music not present in the audio. The generated questions could then ask about some music or sounds not actually present in the audio. We also manually randomized the order of the correct answer within the four choices to ensure a balance.
To verify the potential of such automatically generated training data, we first generated questions for the 40k audios in the AVQA training dataset and name our dataset AVQA-GPT. Using AVQA-GPT, we are able to compare the performance of our models trained on our automatically curated questions vs the original, ground truth human annotated questions. Since the two datasets share the same audio files but have different text question and answers, it would verify the quality of our dataset. In Section IV-B, we show that training on our AVQA-GPT dataset can lead to even better performance than training on the AVQA dataset. Since this approach shows promise, we scaled our method to 182k audio files in the VGGSound training dataset [19] and name this dataset VGGS-GPT. Note that the audio data in AVQA is a subset of the audio data in VGGSound. Due to GPU memory limitations, we filtered the dataset by length of the question and answer pairs, and only used 54k samples from our VGGS-GPT dataset for training. As shown in Section IV-B, training on this extra data could lead to even better performance.
@@
Qwen2.5-Omni-7B Direct Inference [2] - - - - 74.7 68.6 76.0 73.1
@@ IV Experiments
In this section, we report the results on the MMAU [4] and MMAR benchmarks. There are two versions of MMAU. The first is the original version, which we denote as MMAU (Original). All of the previous and concurrent methods have been evaluated on this version. Very recently, an updated version of the benchmark was proposed [^1] https://github.com/Sakshi113/MMAU, which we denote as MMAU (v05.15.25). According to the authors, 25% of the questions and answers have been revised for improved clarity and quality, while 5% of the audio files have been refined to enhance consistency and fidelity. We present the results on MMAU (Original) in Section IV-B and on MMAU (v05.15.25) in Section IV-C. We present the results on MMAR in Section IV-D.
@@ IV-B Main Results on MMAU (Original)
Table I shows the main results on MMAU (Original) [4]. For the baselines, we show the recent methods which achieve State-of-the-Art (SOTA) performance. We refer readers to MMAU [4] for the older baselines [20, 21, 22, 23, 24, 25, 26, 8, 27, 28]. Note that all of the methods were evaluated on the Test-mini split with 1k audio samples, while only a few of them evaluated on the Test-full split with 9k audio samples.
@@
Qwen2-Audio [8] was one of the strongest baselines and several methods are based on it. For example, Zero-Shot-CoT [16] and CoTA [11] try to enhance the model’s reasoning capabilities by encouraging the model to output more of the thinking process. Qwen2.5-Omni [2] is a more recent model which reaches a significantly better average score. Although Kimi-Audio [17] and Audio-Flamingo 2 [13] obtain the best scores on the sound and music categories, Qwen2.5-Omni obtains the best average score of 65.6% (65.9% via our reproduction). SARI [9] adapts Qwen2.5-Omni and improves the average score to 67.1%, which is the current SOTA. However, they calculate the score using an LLM as a judge instead of the usual method based on string matching with the correct answer, therefore their scores might not be directly comparable with the other results. On the larger Test-full set of 9k audios, Audio-Flamingo 2 performs the best on music, while Qwen2.5-Omni obtains the highest scores on the sound, speech, and overall average (68.4%) categories.
@@
Next, we show our Omni-R1 models which fine-tune Qwen2.5-Omni with GRPO. First, compared to the base model, Omni-R1 fine-tuned on AVQA [6] improves the average performance on Test-mini from 65.9% to 68.6% and on Test-full from 68.4% to 70.8%. Next, we show the result of fine-tuning on our proposed AVQA-GPT dataset. It uses the same audio as AVQA but the questions are generated by ChatGPT using audio-based text captions from Qwen-2 Audio [18]. Omni-R1 fine-tuned on AVQA-GPT improves the average performance on Test-mini to 69.9% and on Test-full from to 71.2%. Although AVQA and AVQA-GPT use the same audio files, they use different questions. It verifies the quality of the questions generated by ChatGPT and shows that they are potentially more useful than human generated questions for fine-tuning. Finally, we show the result of fine-tuning on our proposed VGGS-GPT, a larger dataset of questions using the same question generation technique as AVQA-GPT, but scaling to more audio from the VGGSound dataset. It achieves the best average performance of 71.3% on Test-mini and 71.2% on Test-full. Omni-R1 also achieves the highest scores on the sound, music, and speech categories comparing all previous models both on Test-mini and Test-full. It shows that scaling the fine-tuning data can lead to improvements. Compared to the base model, Omni-R1 improved the average scores by 5.4% absolute and 2.8% absolute on Test-mini and Test-full. Finally, compared to SARI which is the most related method and also builds on the same base model, our average Test-mini scores are 4.2% absolute higher.
@@
Gemini 2.0 Flash Direct Inference [29] 61.2 51.0 72.1 81.8 72.5 65.9 70.8 65.6
GPT-4o Audio Direct Inference [30] 53.9 51.0 70.4 63.6 72.5 62.2 75.0 63.5

### v3

@@ I Introduction
Reinforcement Learning (RL) has recently been shown to improve the reasoning capabilities of Large Language Models (LLMs) [guo2025deepseek]. We are motivated by these advancements to improve the capabilities of Audio LLMs - models which take in audio input and text and can perform tasks such as question answering. Building on Qwen2.5-Omni [qwen-omni], a State-of-the-Art (SOTA) multi-modal LLM, we introduce Omni-R1 using a fine-tuning pipeline based on the RL method Group Relative Policy Optimization (GRPO) [shao2024deepseekmath]. Using a simple prompt, our model forgoes complex chain-of-thought or structured reasoning outputs and instead directly outputs answer choices. We use the recent MMAU [sakshi2025mmau] and MMAR [ma2025mmar] benchmarks to test our models. Fine-tuning on the audio and human-annotated questions from the AVQA [yang2022avqa] dataset boosts Qwen2.5-Omni’s average accuracy on MMAU Test-mini from 65.9% to 68.6%, and on the Test-full split from 68.4% to 70.8%.
We also propose to automatically generate audio question answering datasets. We prompt ChatGPT with several audio captions from an audio LLM to generate questions and answer choices for the 40k audios in AVQA and 182k audios in VGGSound. Scaling up the training data with our automatically generated questions results in further gains, achieving new SOTA of 71.3% on Test-mini and 71.2% on Test full. We obtained similar gains on the MMAR [ma2025mmar] benchmark leading to SOTA results.
@@
We plan to release our code, models, and datasets publicly at https://github.com/roudimit/Omni-R1.
@@ II Related Work
Several recent large-scale audio question answering datasets were proposed to test audio LLMs, including MMAU [sakshi2025mmau] and MMAR [ma2025mmar]. The MMAU benchmark [sakshi2025mmau] is a fixed answer choice dataset providing one correct answer and three incorrect answers for each question. It contains questions about sounds, speech, and music of varying difficulty. Some of the questions require external world-based knowledge that is not provided in the questions themselves. MMAR [ma2025mmar] is a more recent benchmark designed to elicit deep reasoning from audio LLMs. It includes questions about sounds, speech, and music, and also mixtures of them (for example, audio natively containing both sounds and speech). The question and answer format is the same as MMAU.
GRPO is a Reinforcement Learning (RL) method proposed for instilling better reasoning capabilities in LLMs [shao2024deepseekmath, guo2025deepseek]. Towards better Audio LLMs, R1-AQA [r1-aqa] proposes to use GRPO by fine-tuning Qwen2-Audio [qwen2-audio] on the AVQA [yang2022avqa] dataset of audio question and answers. This method achieved the previous SOTA on MMAU, and our approach is inspired by them. Using GRPO, we fine-tune Qwen2.5-Omni-7B [qwen-omni], a recent multi-modal LLM proposed to handle both audio and video inputs, on AVQA. Doing so, we were able to achieve the new SOTA on MMAU and also demonstrate strong performance on the MMAR benchmark, achieving SOTA results among open-source models and competitive performance to significantly bigger, close-ended models. Further, we propose to automatically generate audio question and answer training data, which leads to even better performance.
SARI [sari] is a concurrent method which fine-tunes Qwen2.5-Omni using RL. However, their setup is more complex than ours: they fine-tune the model with a schedule of Supervised Fine-Tuning (SFT) and RL, and use both structured reasoning and unstructured reasoning. In contrast, we have a much simpler pipeline: we only fine-tune the model with RL, and we don’t use any explicit reasoning.
Finally, we investigated how GRPO improves the Audio LLMs’ performance by testing them with text-only inputs. While recent work already tested audio LLMs with text-only inputs[sakshi2025mmau, zang2025you], we go a step further and fine-tune the models without audio (just text). We made a surprising discovery that fine-tuning the models with text-only inputs could work almost as well as fine-tuning with audio. It explains that much of the performance improvement from GRPO is due to improving the text-only reasoning capabilities.
@@ III-A Audio LLM and Prompt
We propose Omni-R1 which fine-tunes Qwen2.5-Omni-7B [qwen-omni] using GRPO. During fine-tuning and inference, we use the prompt: <question> Please choose the answer from the following options: <choices>. We chose this prompt since it is simple and the model learns to directly output one of the answer choices. This is especially useful since we have limited GPU memory. Thanks to this simple prompt, we were able to perform full-finetuning on GPUs with only 48GB GPU memory, in contrast to recent works which perform fine-tuning on GPUs with 80GB GPU memory [r1-aqa]. Although recent works [audio-reasoner, r1-aqa, sari] propose more complex prompts with reasoning, our simple prompt was sufficient for our models to achieve the new SOTA on MMAU and MMAR.
@@ III-B Reinforcement Learning
Our training uses Group Relative Policy Optimization (GRPO) [r1-aqa, shao2024deepseekmath], an adaptation of Proximal Policy Optimization (PPO) [schulman2017proximal]. GRPO is designed to mitigate the substantial memory overhead and the challenges of training an accurate per-token value function, common in standard PPO implementations for LLMs. GRPO achieves this by eliminating the explicit value function. Instead, it derives advantage estimates by comparing outputs within a group generated for the same input prompt, effectively using the average reward of sampled responses as a baseline.
@@
—– Human [sakshi2025mmau] 86.3 78.2 82.2 82.2 - - - -
Audio Flamingo 2 Direct Inference [af2] 61.6 74.0 30.9 55.5 65.1 72.9 40.3 59.4
Phi-4 Multi-Modal Direct Inference [phi-4] 61.0 52.9 52.8 55.6 - - - -
GPT-4o Voice Model Direct Inference [gpt-4o] 63.4 60.8 53.2 59.1 - - - -
Qwen2-Audio-7B-Instruct Direct Inference [qwen2-audio] 55.0 51.0 42.0 49.2 45.9 53.3 45.9 52.5
Qwen2-Audio-7B-Instruct Reproduced [qwen2-audio] 61.6 54.5 42.0 52.7 - - - -
Qwen2-Audio-7B-Instruct Zero-Shot-CoT [audio-cot] 61.9 56.3 55.3 57.8 - - - -
Qwen2-Audio-7B-Instruct CoTA [audio-reasoner] 60.1 64.3 60.7 61.7 - - - -
Qwen2-Audio-7B-Instruct R1-AQA [r1-aqa] 68.8 64.3 63.7 65.6 69.8 61.4 62.7 64.4
Kimi-Audio Direct Inference [kimi] 73.3 61.7 60.7 65.2 - - - -
Qwen2.5-Omni-7B Direct Inference [qwen-omni] 67.9 69.2 59.8 65.6 - - - -
@@
Qwen2.5-Omni-7B SARI [sari] 72.8 67.2 61.3 67.1 - - - -
@@
$\footnotesize\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{\mathcal{D}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg\{\min\Bigg[\rho_{i,t}\hat{A}_{i,t},\\ &\qquad\left.\operatorname{clip}\left(\rho_{i,t},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta{\mathbb{D}_{KL}}_{i,t}\left[\pi_{\theta}||\pi_{ref}\right]\Bigg\}\Bigg],\end{split}$ (3)
where $\rho_{i,t}=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}$ is the probability ratio between the current policy $\pi_{\theta}$ and the policy $\pi_{\theta_{old}}$ used for sampling and $\mathcal{D}=[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)]$ denote the sampling distribution over questions and their corresponding outputs. The $\operatorname{clip}(x,min\_val,max\_val)$ function limits the value of $x$ to be within the range [min_val,max_val][min\_val,max\_val]. The hyperparameter $\epsilon$ defines the PPO clipping range. The per-token KL divergence ${\mathbb{D}_{KL}}_{i,t}\left[\pi_{\theta}||\pi_{ref}\right]$ between the current policy $\pi_{\theta}$ and a reference policy $\pi_{ref}$ is weighted by the hyperparameter $\beta$:
@@ III-C Automatic Q/A Dataset Generation with ChatGPT
We propose to generate audio question answer datasets automatically. Given an audio dataset, we first acquire audio-based text captions from an audio LLM. Specifically, we use the captions generated by Qwen-2 Audio [qwen2-audio] in the AudioSetCaps dataset [bai2024audiosetcaps], which is a compilation of captions generated for many open-source audio datasets. We use multiple captions which describe the overall acoustic scene, the speech present in the audio, and the music present in the audio. Next, we use the captions and prompt ChatGPT to generate questions and answers using the provided captions. We instruct the model to generate a correct answer and three other answers which are similar but clearly incorrect (prompt shown in Figure 1). We reviewed the generated questions and answers and found the quality to be reasonable, however, we noticed that the original captions feature hallucinations especially about music not present in the audio. The generated questions could then ask about some music or sounds not actually present in the audio. We also manually randomized the order of the correct answer within the four choices to ensure a balance.
To verify the potential of such automatically generated training data, we first generated questions for the 40k audios in the AVQA training dataset and name our dataset AVQA-GPT. Using AVQA-GPT, we are able to compare the performance of our models trained on our automatically curated questions vs the original, ground truth human annotated questions. Since the two datasets share the same audio files but have different text question and answers, it would verify the quality of our dataset. In Section IV-B, we show that training on our AVQA-GPT dataset can lead to even better performance than training on the AVQA dataset. Since this approach shows promise, we scaled our method to 182k audio files in the VGGSound training dataset [chen2020vggsound] and name this dataset VGGS-GPT. Note that the audio data in AVQA is a subset of the audio data in VGGSound. Due to GPU memory limitations, we filtered the dataset by length of the question and answer pairs, and only used 54k samples from our VGGS-GPT dataset for training. As shown in Section IV-B, training on this extra data could lead to even better performance.
@@
Qwen2.5-Omni-7B Direct Inference [qwen-omni] - - - - 74.7 68.6 76.0 73.1
@@ IV Experiments
In this section, we report the results on the MMAU [sakshi2025mmau] and MMAR benchmarks. There are two versions of MMAU. The first is the original version, which we denote as MMAU (Original). All of the previous and concurrent methods have been evaluated on this version. Very recently, an updated version of the benchmark was proposed [^1] https://github.com/Sakshi113/MMAU, which we denote as MMAU (v05.15.25). According to the authors, 25% of the questions and answers have been revised for improved clarity and quality, while 5% of the audio files have been refined to enhance consistency and fidelity. We present the results on MMAU (Original) in Section IV-B and on MMAU (v05.15.25) in Section IV-C. We present the results on MMAR in Section IV-D.
@@ IV-B Main Results on MMAU (Original)
Table I shows the main results on MMAU (Original) [sakshi2025mmau]. For the baselines, we show the recent methods which achieve State-of-the-Art (SOTA) performance. We refer readers to MMAU [sakshi2025mmau] for the older baselines [deshmukh2023pengi, gong2024listen, gong2023joint, deng2024musilingo, liu2024music, liu2023m, ghosh2024gama, qwen2-audio, tangsalmonn, qwenaudio]. Note that all of the methods were evaluated on the Test-mini split with 1k audio samples, while only a few of them evaluated on the Test-full split with 9k audio samples.
@@
Qwen2-Audio [qwen2-audio] was one of the strongest baselines and several methods are based on it. For example, Zero-Shot-CoT [audio-cot] and CoTA [audio-reasoner] try to enhance the model’s reasoning capabilities by encouraging the model to output more of the thinking process. Qwen2.5-Omni [qwen-omni] is a more recent model which reaches a significantly better average score. Although Kimi-Audio [kimi] and Audio-Flamingo 2 [af2] obtain the best scores on the sound and music categories, Qwen2.5-Omni obtains the best average score of 65.6% (65.9% via our reproduction). SARI [sari] adapts Qwen2.5-Omni and improves the average score to 67.1%, which is the current SOTA. However, they calculate the score using an LLM as a judge instead of the usual method based on string matching with the correct answer, therefore their scores might not be directly comparable with the other results. On the larger Test-full set of 9k audios, Audio-Flamingo 2 performs the best on music, while Qwen2.5-Omni obtains the highest scores on the sound, speech, and overall average (68.4%) categories.
@@
Next, we show our Omni-R1 models which fine-tune Qwen2.5-Omni with GRPO. First, compared to the base model, Omni-R1 fine-tuned on AVQA [yang2022avqa] improves the average performance on Test-mini from 65.9% to 68.6% and on Test-full from 68.4% to 70.8%. Next, we show the result of fine-tuning on our proposed AVQA-GPT dataset. It uses the same audio as AVQA but the questions are generated by ChatGPT using audio-based text captions from Qwen-2 Audio [bai2024audiosetcaps]. Omni-R1 fine-tuned on AVQA-GPT improves the average performance on Test-mini to 69.9% and on Test-full from to 71.2%. Although AVQA and AVQA-GPT use the same audio files, they use different questions. It verifies the quality of the questions generated by ChatGPT and shows that they are potentially more useful than human generated questions for fine-tuning. Finally, we show the result of fine-tuning on our proposed VGGS-GPT, a larger dataset of questions using the same question generation technique as AVQA-GPT, but scaling to more audio from the VGGSound dataset. It achieves the best average performance of 71.3% on Test-mini and 71.2% on Test-full. Omni-R1 also achieves the highest scores on the sound, music, and speech categories comparing all previous models both on Test-mini and Test-full. It shows that scaling the fine-tuning data can lead to improvements. Compared to the base model, Omni-R1 improved the average scores by 5.4% absolute and 2.8% absolute on Test-mini and Test-full. Finally, compared to SARI which is the most related method and also builds on the same base model, our average Test-mini scores are 4.2% absolute higher.
@@
Gemini 2.0 Flash Direct Inference [google_gemini_2_flash_2025] 61.2 51.0 72.1 81.8 72.5 65.9 70.8 65.6
GPT-4o Audio Direct Inference [gpt-4o-openai] 53.9 51.0 70.4 63.6 72.5 62.2 75.0 63.5